The rate costant K1 of a reaction is found to be double of that of rate constant K2 of another reaction .What is the relationship between corresponding activation energies of the to reactions .The Arrhenius constant is same for the reactions.


(A)E2 = E1 - ln 2 /RT


(B)E2=E1 ln 2/RT


(C)E2 + E1 = ln 2/RT


(D) E2 =E1 + RTln 2


Ans:(C)


Thanks in advance.

Grade:12

1 Answers

ANMOL SARIN
14 years ago

Given is that

k1  = 2*k2               .................(1)

Let A be the arrhenius constant for both the reactions . Then using arrhenius equation,       

k1=A*exp(-E1/RT)    ..............(2)

k2=A*exp( -E2/RT)  ...............(3)

[ where exp stands for 'e' raised to the power of ]

Divide (2) by (3)

k1/k2=exp(E2-E1)/RT

OR exp(E2-E1)/RT= 2   [ from equation (1)  ]

Taking log both sides to the base e ,

 

(E2-E1)/RT = ln2

E2 = E1 + RTln2

Ans is (D ) & not (C)

If you watch all four options, equations given in (B) & (C) are dimensionally incorrect . So they should be ruled out at once.

Calculations show answer to be (D) as above.